[figure: 9.7 cm]

MAGNETIC VISCOSITY IN VANADIUM CHROMICYANIDES

M. Sendek, V. Kavečanský, M. Lukáčová, M. Maryško^{b}, Z. Mitróová,A. Zentko

Institute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53

Košice, Slovak Republic bInstitute of Physics ASCR, Na Slovance 2 Prague, Czech Republic

We have synthesized V [Cr(CN)_{6}] · nH_{2}O and V O[Cr(CN)_{6}] · nH_{2}O molecular based magnets belonging to the Prussian blue family of compounds. Previous results [1] demonstrate that the transition – metal hexacyano complexes are promising components for the construction of molecular – based magnets high – T_{c} magnets. In this paper we report on the synthesis, characterization and magnetic viscosity in V O[Cr(CN)_{6}]·nH_{2}O and V [Cr(CN)_{6}] · nH_{2}O compounds. Obtained results display a logarithmic behavior of the magnetization versus time dependence in accordance with Street-Woolley theory based on activation energy concepts [2]. [1] S. Ferlay, T. Mallah, R. Quahes, P. Veilleth,

M. Verdaguer: Nature 378 (1995) 701.

[2] R. Street, J.C. Wooley: Proc.Phys. Soc., A62 (1949) 562.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Sendek

Address for correspondence : Institute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovak Republic

Email address : sendek@saske.sk